OpenStudy (anonymous):

Calculus Find critical numbers of the function f(x) 4x/ x^2+1

OpenStudy (anonymous):

\[f(x)\frac{ 4x }{ x^2+1 }\] \[f'(x)=\frac{4(x^2+1)-[4x(2x)] }{ (x^2+1)^2 }\]

OpenStudy (shamil98):

Looks good so far. You can cancel out from the numerator and denominator i think.

OpenStudy (shamil98):

(x^2 + 1)^2 = 0 x^2 + 1 = 0 x^2 = 1 \[x = \pm 1\] at + or - 1 your points cannot exist there

OpenStudy (tkhunny):

\(x^{2} = -1\) There are NO Real Solutions to this.

OpenStudy (shamil98):

oh oops, brainfarted there

OpenStudy (anonymous):

Do you mean like this \[f'(x)={ \frac{ 4x^2+4-8x^2 }{ (x^2+1)^2} }\] f'(x)=\[\frac{ -4x^2+4 }{ (x^2+1)^2 }\]

hartnn (hartnn):

ok thats correct by far to find critical points, you equate f'(x) = 0 so your numerator -4x^2+4 = 0 you'll get 2 real (integer) values of x from this

OpenStudy (tkhunny):

It looks really tempting to simplify this expression a bit. Don't do it. \(\dfrac{-4x^{2}+4}{x^{2}+1} = -4\cdot\dfrac{x^{2} - 1}{x^{2}+1}\) Nothing to simplify, there.